OpenStudy (anonymous):

Assuming x > 0, which of these expressions is equivalent to

OpenStudy (anonymous):

\[11\sqrt{245x ^{3}}+9\sqrt{45x ^{3}}\]

OpenStudy (anonymous):

anwser: \[5x \sqrt{104x}-20\sqrt{290x ^{6}}-20x \sqrt{290x}-104x \sqrt{5x}\]